OpenStudy (anonymous):

Does the sum of a convergent and divergent series always convergeee? My textbook says nothing about this

OpenStudy (perl):

the sum of a convergent and divergent series never converges

OpenStudy (perl):

we can prove it

OpenStudy (perl):

Suppose you are given a convergent series a_n and a divergent series b_n. If the sum exists, call it c. $$ \Large { \sum a_n + \sum b_n = c \\ \sum b_n = c - \sum a_n } $$

OpenStudy (perl):

now you have a contradiction

OpenStudy (perl):

we can also write it like this. Suppose that Convergent series + divergent series = k , for some constant k subtract both sides the convergent series divergent series = k - convergent series but the right side is a finite number, since its k - a and the left side is divergent

OpenStudy (anonymous):

Ah, thank yoooou.

OpenStudy (perl):

This might be clearer. $$ \Large { \sf suppose~ that \\ \sum a_n = k, \sum b_n = does~ not~ exist \\ If ~~\sum a_n + \sum b_n = C \\~\\ then \\~\\ \sum b_n = C - \sum a_n \\ \sum b_n = C - A \\~\\ then \\~\\ \sum b_n~ does ~exist \\ ~\\ contradiction! \\ } $$

OpenStudy (anonymous):

Thank you for the fast response. =)